The kinetic energy needed to project a body of mass m from the earth surface (radius R) to infinity is
(a) mgR/2
(b) 2 mg /R
(c) mgR
(d) mgR/4

Grade:6

3 Answers

Deepak Patra
9 years ago

(c) K.E. = ½ mv2e ­where ve = escape velocity = v2gR

? K.E = ½ m x 2gR = mgR
Abhishek Dimri
6 years ago
The answer is (c)We know that K.E. 1/2 mV^2e. ..........(1)Where Ve = escape velocityVe = √(2gR)V^2e = 2gRPutting the value of V^2e in eqn (1) we have K.E. = 1/2 m2gRK.E. = mgR
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Since only conservative force(Gravitational force acts), net mechanical energy is conserved.
Hence, Ki + Ui = Kf + Uf
Hence, Ki + ( – GMm/R) = 0 + 0
Hence, Ki = GMm/R = GM/R x m = GM/R2 x m x R = mgR
 
Hope it helps,
Thanks and regards,
Kushagra